OpenStudy (anonymous):

The ordered pairs (1,81),(2,100),(3,121),(4,144), and (5, 169) represent a function. What is the rule that represents this function?

OpenStudy (anonymous):

81= (1+8)^2 100=(2+8)^2 121=(3+8)^2 144=(4+8)^2 169=(5+8)^2 Got it?

OpenStudy (anonymous):

\[ f(x)=(x+8)^2 \]